What is the percent mass of oxygen in 87 g of hydrogen peroxide
1
Expert's answer
2020-11-09T13:53:59-0500

Solution:

Molar mass of H2O2 = 34.01 g/mol

Mass of O in 1 mol H2O2 = 2×16 = 32 g

%O = 32g O / 34.01g H2O2 × 100% = 94.09 %

Mass O = 32g O / 34.01g H2O2 × 87g H2O2 = 81.86 g

Answer: the percent mass of oxygen is 94.09% or 81.86 grams of O.